OpenStudy (anonymous):

how do i find the tenth derivatve of y=1/(2x-1)?

OpenStudy (anonymous):

lol your kidding right?

OpenStudy (anonymous):

maybe if you take successive derivatives you will see a pattern why do you need the tenth derivative?

OpenStudy (anonymous):

\[y=(2x-1)^{-1}\] usually it is not helpful to use exponential notation but it probably is here \[y'=-2(x-1)^{-2}\] \[y''=4(2x-1)^{-3}\] \[y'''=-12(2x-1)^{-4}\] hmmm looks promising

OpenStudy (anonymous):

more promising if i had not made a typo \[y'=-2(2x-1)^{-1}\]

OpenStudy (anonymous):

need something like \((-1)^n\) to make in alternate and also \(n!\) or maybe \(2\times n!\)